If it requires 3.0 J of work to stretch a particular spring by 2.1 cm from its equilibrium length, how much more work will be required to stretch it an additional 3.7 cm ?

Express your answer using two significant figures.
Answer #1

given

x = 2.1 cm

= 0.021 m

U = 3 J

k = spring constant

work done in the spring is U = 1/2 k x2

3 = 1/2 x k x 0.0212

k = 3 / 0.5 x 0.0212

k = 13605.442 N/m

to stretch to 2.1 + 3.7 = 5.8 cm

= 0.058 m

then the work done is U' = 1/2 k x2

= 0.5 x 13605.442 x 0.0582

U ' = 22.8843 J

to stretch 3.7 cm work need is = 22.8843 - 3 = 19.8843 J

then the more work will be required to stretch it an additional 3.7 cm is 19.8843 J